# The radius of statistical efficiency

### v1

@@ D.2 Proof of Theorem 4
$\nabla^{2}f(M_{\star})[\Delta,\Delta]$ $\geq\frac{2}{\kappa\left(\Sigma_{1}\right)\kappa\left(\Sigma_{2}\right)+1}\left(\langle\beta_{1\star},\Sigma_{1}\beta_{1\star}\rangle\langle u,\Sigma_{2}u\rangle+\langle\beta_{2\star},\Sigma_{2}\beta_{2\star}\rangle\langle a\beta_{1\star}+v,\Sigma_{1}(a\beta_{1\star}+v)\rangle\right)$ (120)
@@
$\geq\frac{2}{\kappa\left(\Sigma_{1}\right)\kappa\left(\Sigma_{2}\right)+1}\left(\langle\beta_{1\star},\Sigma_{1}\beta_{1\star}\rangle\lambda_{\min}\left(\Sigma_{2}\right)\|u\|^{2}+\langle\beta_{2\star},\Sigma_{2}\beta_{2\star}\rangle\lambda_{\min}\left(\Sigma_{1}\right)\|a\beta_{1\star}+v\|^{2}\right)$ (121)
@@
$\geq\frac{2}{\kappa\left(\Sigma_{1}\right)\kappa\left(\Sigma_{2}\right)+1}\min\left\{\langle\beta_{1\star},\Sigma_{1}\beta_{1\star}\rangle\lambda_{\min}\left(\Sigma_{2}\right),\langle\beta_{2\star},\Sigma_{2}\beta_{2\star}\rangle\lambda_{\min}\left(\Sigma_{1}\right)\right\}$ (122)
where the last holds since $\|u\|^{2}+\|a\beta_{1\star}+v\|^{2}=1.$ This establishes the lower bound in (32).
Next we establish the converse. For any $\Delta=a\beta_{1\star}\beta_{2\star}^{\top}+u\beta_{2\star}^{\top}+\beta_{1\star}v^{\top}$, from (100) we have
@@
$\nabla^{2}f(M_{\star})[\Delta,\Delta]=\operatorname{\mathbb{E}}[a\big{\langle}x_{1},\beta_{1\star}\big{\rangle}\big{\langle}x_{2},\beta_{2\star}\big{\rangle}+\big{\langle}x_{1},u\big{\rangle}\big{\langle}x_{2},\beta_{2\star}\big{\rangle}+\big{\langle}x_{1},\beta_{1\star}\big{\rangle}\big{\langle}x_{2},v\big{\rangle}]^{2}.$ (123)
Note the equality $\|\Delta\|^{2}_{F}=a^{2}\|\beta_{1\star}\|_{2}^{2}\|\beta_{2\star}\|_{2}^{2}+\|\beta_{1\star}\|^{2}\|v\|^{2}+\|\beta_{2\star}\|^{2}\|u\|^{2}$. Now setting $v=0$, equation (123) becomes
@@
$\nabla^{2}f(M_{\star})[\Delta,\Delta]=\operatorname{\mathbb{E}}[\big{\langle}x_{1},a\beta_{1\star}+u\big{\rangle}\big{\langle}x_{2},\beta_{2\star}\big{\rangle}]^{2}=\langle\Sigma_{1}(a\beta_{1\star}+u),a\beta_{1\star}+u\rangle\langle\Sigma_{2}\beta_{2\star},\beta_{2\star}\rangle.$
Note the equality $\|a\beta_{1\star}+u\|^{2}=a^{2}\|\beta_{1\star}\|^{2}+\|u\|^{2}=\frac{\|\Delta\|^{2}}{\|\beta_{2\star}\|^{2}}$. Therefore we deduce
@@
$\min_{\Delta:\|\Delta\|=1}\nabla^{2}f(M_{\star})[\Delta,\Delta]\leq\min_{w:\|w\|^{2}=1}\langle\Sigma_{1}w,w\rangle\langle\Sigma_{2}\beta_{2\star},\beta_{2\star}\rangle/\|\beta_{2\star}\|^{2}=\lambda_{\min}(\Sigma_{1})\cdot\langle\Sigma_{2}\beta_{2\star},\beta_{2\star}\rangle/\|\beta_{2\star}\|^{2}.$
A symmetric argument shows
$\min_{\Delta:\|\Delta\|=1}\nabla^{2}f(M_{\star})[\Delta,\Delta]\leq\lambda_{\min}(\Sigma_{2})\cdot\langle\Sigma_{1}\beta_{1\star},\beta_{1\star}\rangle/\|\beta_{1\star}\|^{2}.$
@@
In particular, if $\Sigma_{1}$ or $\Sigma_{2}$ are singular then $\mu\times\nu$ lies in $\mathcal{E}$, as claimed. This completes the proof.
@@
Consider any tangent vector $\Delta={\beta}_{\star}v^{\top}+v{\beta}_{\star}^{\top}$ with $v\in\mathbf{R}^{d}$. Then we may vectorize $\Delta$ as follows:
$\operatorname{vec}(\Delta)=\operatorname{vec}({\beta}_{\star}v^{\top}I)+\operatorname{vec}(Iv{\beta}_{\star}^{\top})=(I\otimes{\beta}_{\star})v+({\beta}_{\star}\otimes I)v=\Phi_{\beta}v.$
Therefore, we compute
$\nabla f(M_{\star})[\Delta,\Delta]=\operatorname{vec}(\Delta)^{\top}\Sigma_{\mu}{\operatorname{vec}}(\Delta)=v^{\top}(\Phi_{\beta_{\star}}^{\top}\Sigma_{\mu}\Phi_{\beta_{\star}})v,$ (124)
Minimizing the expression (124) in $\Delta$ with $\|\Delta\|_{F}=1$ yields the guarantee (37).
@@
Characterization of well-posedness. To simplify notation, we will relabel ${\beta}_{\star}$ to $\beta$. Observe for any measure $(p_{ij})\in\mathcal{Q}$, equation (35) yields
$\nabla^{2}f(M_{\star})[\Delta_{v},\Delta_{v}]=\operatorname{\mathbb{E}}\langle\Delta_{v},X\rangle^{2}=\sum_{(i,j)\in E}(v_{i}\beta_{j}+v_{j}\beta_{i})^{2}p_{ij},$ (125)
where $\Delta_{v}:=\beta v^{\top}+v\beta^{\top}$ is any tangent vector to $M_{\star}$ at $\mathcal{M}$. Recall moreover that $\Delta_{v}$ is nonzero if and only if $v$ is nonzero.
($\implies$) We prove the contrapositive. Thus, suppose that Assumption 1 does not hold. We will consider two cases separately.
Case 1. Assume that there exists $\bar{i}\in V^{0}$. Thus for any edge $(\bar{i},j)\in E$ equality $\beta_{j}=0$ holds. Then setting $v=e_{\bar{i}}$, we deduce
$\nabla^{2}f(M_{\star})[\Delta_{v},\Delta_{v}]=\sum_{(i,j)\in E}(v_{i}\beta_{j}+v_{j}\beta_{i})^{2}p_{ij}=\sum_{j:(\bar{i},j)\in E}\beta_{j}^{2}p_{\bar{i}j}=0.$
We conclude that $\mu$ lies in $\mathcal{E}^{{\rm mc}}$, as claimed.
Case 2. Assume that one of the components of $G^{*}$ is bipartite and the set $V^{0}$ is empty. Without loss of generality, we may assume that $G^{*}$ is connected, since otherwise we can restrict the following argument to any connected component. Thus, there exists a partition $V^{*}=I\cup J$ with all the edges $(i,j)\in E^{*}$ satisfying $i\in I$ and $j\in J$. Define the vector
$v_{i}=\begin{cases}\beta_{i}&\text{if }i\in I,\\ -\beta_{j}&\text{if }j\in J,\\ 0&\text{otherwise.}\end{cases}$
Using (125), we obtain
$\nabla^{2}f(M_{\star})[\Delta_{v},\Delta_{v}]=\sum_{(i,j)\in E^{*}}(v_{i}\beta_{j}+v_{j}\beta_{i})^{2}p_{ij}=\sum_{(i,j)\in E^{*}}(\beta_{i}\beta_{j}-\beta_{j}\beta_{i})^{2}p_{ij}=0.$ (126)
where we used that $v$ is supported on $V^{*}$ and that $V^{0}$ is empty. Thus, $\mu$ lies in $\mathcal{E}^{{\rm mc}}$, as claimed.
@@
$v_{i}\beta_{j}+v_{j}\beta_{i}=0\qquad\forall(i,j)\in E.$ (127)
Without loss of generality suppose that $G^{*}$ is connected, otherwise, we can repeat the argument for each connected component. Since $G^{*}$ is non-bipartite, it must contain an odd-size cycle $i_{1}\rightarrow i_{2}\rightarrow\dots\rightarrow i_{k}\rightarrow i_{1}$. Consider the expansion
$2\beta_{i_{1}}v_{i_{k}}=\left(\beta_{i_{1}}v_{i_{k}}+\beta_{i_{k}}v_{i_{1}}\right)-\frac{\beta_{i_{k}}}{\beta_{i_{2}}}\left(\beta_{i_{2}}v_{i_{1}}+\beta_{i_{1}}v_{i_{2}}\right)+\sum_{j=2}^{k-1}(-1)^{j}\frac{\beta_{i_{1}}\beta_{i_{k}}}{\beta_{i_{j}}\beta_{i_{j+1}}}\left(\beta_{i_{j+1}}v_{i_{j}}+\beta_{i_{j}}v_{i_{j+1}}\right)=0,$
where the last equality follows since each term in the parenthesis is zero. Since $\beta_{i_{1}}>0$, we deduce $v_{i_{k}}=0$. Next, observe from (127) that for any neighbor $j$ of $i_{k}$, i.e., satisfying $(i_{k},j)\in E^{*}$, we have that $v_{j}=\beta_{j}v_{i_{k}}/\beta_{i_{k}}=0.$ Repeating the argument, we deduce $v_{j}=0$ for all $v\in V^{*}$. Next, consider any vertex $i\notin V^{*}$. Then since $V^{0}$ is empty, there exists some $j\in V^{*}$ with $(i,j)\in E$. Using (127) again, we conclude $v_{i}\beta_{j}=0$. Taking into account $\beta_{j}>0$, we deduce $v_{i}=0$. Thus $v$ is identically zero.
@@
We have now proved that a measure $\mu$ lies in $\mathcal{E}^{{\rm mc}}$ if and only if its support $\operatorname{supp}(\mu)$ lies in $\Omega_{{\beta}_{\star}}$. Given any set of indices $A\subseteq[d]\times[d]$ define $\mathsf{P}_{A}\colon{\mathbf{R}}^{d\times d}\rightarrow{\mathbf{R}}^{d\times d}$ to be the orthogonal projection onto the entries indexed by $A$. Fix now a set of entries $A\subset\operatorname{supp}(P)$ such that $A\in\Omega_{M_{\star}}.$ Clearly, the pushforward measure $(\mathsf{P}_{A})_{\#}\mu$ lies $\mathcal{E}^{\text{mc}}$, and we compute
$\min_{\nu\in\mathcal{E}^{\text{mc}}}W_{2}^{2}(\mu,\nu)$ $\leq W_{2}^{2}\big{(}\mu,(\mathsf{P}_{A})_{\#}\mu\big{)}=\sum_{ij\in\operatorname{supp}(P)\setminus A}p_{ij}.$
Taking the minimum over $A$ yields the inequality $\leq$ in (40).
Conversely, fix a measure $\nu\in\mathcal{E}^{\text{mc}}$ and let $A$ be the indices that are observed with positive probability according to $\nu$. Thus, the support of $\nu$ is contained in the subspace $\operatorname{range}(\mathsf{P}_{A})$. Then by Lemma A.1 and (42), we have
$W_{2}^{2}(\mu,\nu)\geq W_{2}^{2}(\mu,(\mathsf{P}_{A})_{\#}\mu)=\sum_{ij\in\operatorname{supp}(P)\setminus A}p_{ij}.$
Taking the minimum over all entries $A\subset\operatorname{supp}(P)$ such that $A\in\Omega_{M_{\star}}$ completes the proof of the reverse inequality $\geq$ in (40).
@@
We reduce from the MAXCUT problem. Assume that we have a polynomial time algorithm Alg$(M_{\star},(p_{ij}))$ to compute (40). Given an instance of MAXCUT $G=(V,E)$, we split the graph into its connected components $G_{1}=(V_{1},E_{1}),\dots,G_{k}=(V_{k},E_{k})$. For each $\ell\leq k$, we define an instance of matrix completion with $M_{\star}=\mathbf{1}\mathbf{1}^{\top}\in{\mathbf{R}}^{|G_{\ell}|\times|G_{\ell}|}$ — with a slight abuse of notation we use $M_{\star}$ for all $k$ problems — and set the distribution ${P}^{(\ell)}$ to $p_{ij}^{(\ell)}=\frac{1}{|E_{\ell}|}$ if $i,j\in E_{\ell}$, and $p_{ij}^{(\ell)}=0$, otherwise. Since the entries of $M_{\star}$ are strictly positive and $G_{\ell}$ is connected, the output of Alg$(M_{\star},{P}^{(\ell)})$ times $|E_{\ell}|$ is equal to the minimum number of edges one needs to remove from $G_{\ell}$ to make it bipartite. Thus, $|E_{\ell}|\left(1-\texttt{ALG}\left(M_{\star},{P}^{(\ell)}\right)\right)$ is equal to the number of edges of the largest bipartite graph one can construct via edge deletion, which is readily seen to be equal to $\texttt{MAXCUT}(G_{\ell}).$ Thus summing along connected components yields
$\texttt{MAXCUT}(G)=|E|-\sum_{\ell\leq k}|E_{\ell}|\cdot\texttt{ALG}\Big{(}M_{\star},{P}^{(\ell)}\Big{)};$
completing the proof.

### v2

@@ D.2 Proof of Theorem 4
$\nabla^{2}f(M_{\star})[\Delta,\Delta]$ $\geq\frac{2}{\kappa(\Sigma_{1})\kappa(\Sigma_{2})+1}\Big(\langle\beta_{2\star},\Sigma_{2}\beta_{2\star}\rangle\langle w,\Sigma_{1}w\rangle+\langle\beta_{1\star},\Sigma_{1}\beta_{1\star}\rangle\langle v,\Sigma_{2}v\rangle\Big)$ (124)
@@
$\geq\frac{2}{\kappa(\Sigma_{1})\kappa(\Sigma_{2})+1}\Big(\langle\beta_{2\star},\Sigma_{2}\beta_{2\star}\rangle\lambda_{\min}(\Sigma_{1})\|w\|^{2}_{2}+\langle\beta_{1\star},\Sigma_{1}\beta_{1\star}\rangle\lambda_{\min}(\Sigma_{2})\|v\|^{2}_{2}\Big)$ (125)
$\geq\frac{2\cdot{\min}\{\gamma_{2}\lambda_{\min}(\Sigma_{1}),\gamma_{1}\lambda_{\min}(\Sigma_{2})\}}{\kappa(\Sigma_{1})\kappa(\Sigma_{2})+1}\|\Delta\|_{F}^{2},$ (126)
@@
where $\gamma_{i}:=\left\langle\Sigma_{i}\frac{\beta_{i\star}}{\|\beta_{i\star}\|_{2}},\frac{\beta_{i\star}}{\|\beta_{i\star}\|_{2}}\right\rangle$ for $i=1,2$. This establishes the lower bound in (34).
Next, we establish the upper bound in (34). Observe that from (105), we have
@@
$\nabla^{2}f(M_{\star})[\Delta,\Delta]={\operatorname{\mathbb{E}}}\big[\langle x_{1},w\rangle\langle x_{2},\beta_{2\star}\rangle+\langle x_{1},\beta_{1\star}\rangle\langle x_{2},v\rangle\big]^{2}.$ (127)
If $v=0$ and $w\neq 0$, then (127) becomes
@@
$\nabla^{2}f(M_{\star})[\Delta,\Delta]={\operatorname{\mathbb{E}}}\big[\langle x_{1},w\rangle\langle x_{2},\beta_{2\star}\rangle\big]^{2}=\langle\Sigma_{1}w,w\rangle\langle\Sigma_{2}\beta_{2\star},\beta_{2\star}\rangle=\gamma_{2}\Big\langle\Sigma_{1}\tfrac{w}{\|w\|_{2}},\tfrac{w}{\|w\|_{2}}\Big\rangle\|\Delta\|^{2}_{F}$ (128)
and hence minimizing over all such $\Delta$ yields
$\lambda_{\min}(\nabla^{2}f(M_{\star}))\leq\gamma_{2}\lambda_{\min}(\Sigma_{1}).$ (129)
Similarly, if $w=0$ and $v\neq 0$, then (127) becomes
@@
$\nabla^{2}f(M_{\star})[\Delta,\Delta]={\operatorname{\mathbb{E}}}\big[\langle x_{1},\beta_{1\star}\rangle\langle x_{2},v\rangle\big]^{2}=\langle\Sigma_{1}\beta_{1\star},\beta_{1\star}\rangle\langle\Sigma_{2}v,v\rangle=\gamma_{1}\Big\langle\Sigma_{2}\tfrac{v}{\|v\|_{2}},\tfrac{v}{\|v\|_{2}}\Big\rangle\|\Delta\|^{2}_{F}$ (130)
and hence minimizing over all such $\Delta$ yields
$\lambda_{\min}(\nabla^{2}f(M_{\star}))\leq\gamma_{1}\lambda_{\min}(\Sigma_{2}).$ (131)
@@
Thus, $\lambda_{\min}(\nabla^{2}f(M_{\star}))\leq{\min}\{\gamma_{2}\lambda_{\min}(\Sigma_{1}),\gamma_{1}\lambda_{\min}(\Sigma_{2})\}$ and the upper bound in (34) is established. In particular, if either $\Sigma_{1}$ or $\Sigma_{2}$ are singular, then $\mu\times\nu$ lies in $\mathcal{E}$, as claimed. This completes the proof.
@@
Consider any tangent vector $\Delta={\beta}_{\star}v^{\top}+v{\beta}_{\star}^{\top}\in{\mathcal{T}}$, where $v\in\mathbf{R}^{d}$. We may vectorize $\Delta$ as follows:
$\operatorname{vec}(\Delta)=\operatorname{vec}({\beta}_{\star}v^{\top}I)+\operatorname{vec}(Iv{\beta}_{\star}^{\top})=(I\otimes{\beta}_{\star})v+({\beta}_{\star}\otimes I)v=\Phi_{{\beta}_{\star}}v.$
Therefore, from (37) we compute
$\nabla^{2}f(M_{\star})[\Delta,\Delta]=\operatorname{vec}(\Delta)^{\top}\Sigma_{\mu}{\operatorname{vec}}(\Delta)=v^{\top}(\Phi_{\beta_{\star}}^{\top}\Sigma_{\mu}\Phi_{\beta_{\star}})v.$ (132)
Assuming $v\neq 0$, we recall $\|\Delta\|_{F}^{2}/\|{\beta}_{\star}\|^{2}_{2}\|v\|^{2}_{2}\in[2,4]$ and hence (132) implies
$\frac{\big\langle(\Phi_{\beta_{\star}}^{\top}\Sigma_{\mu}\Phi_{\beta_{\star}})v,v\big\rangle}{4\|{\beta}_{\star}\|_{2}^{2}\|v\|_{2}^{2}}\leq\frac{\nabla^{2}f(M_{\star})[\Delta,\Delta]}{\|\Delta\|_{F}^{2}}\leq\frac{\big\langle(\Phi_{\beta_{\star}}^{\top}\Sigma_{\mu}\Phi_{\beta_{\star}})v,v\big\rangle}{2\|{\beta}_{\star}\|_{2}^{2}\|v\|_{2}^{2}}.$ (133)
Minimizing the ratios in (133) over all $v\in\mathbf{R}^{d}\setminus\{0\}$ yields the estimate (39).
@@
$\nabla^{2}f(M_{\star})[\Delta,\Delta]=\operatorname{\mathbb{E}}\langle X,\Delta\rangle^{2}=\sum_{i,j:\,p_{ij}>0}(v_{i}\beta_{j}+v_{j}\beta_{i})^{2}p_{ij}.$ (134)
$(\!\implies\!)$ We prove the contrapositive. Thus, suppose $G$ does not satisfy Assumption 1. We consider two cases separately.
Case 1. Suppose $G^{*}$ has a bipartite connected component $G_{\ell}^{*}=(V^{*}_{\ell},E^{*}_{\ell})$, so that there exists a disjoint decomposition $V^{*}_{\ell}=I\sqcup J$ of $V^{*}_{\ell}\subset\operatorname{supp}(\beta)$ such that every edge in $E^{*}_{\ell}$ is of the form $\{i,j\}$ for some $(i,j)\in I\times J$. Then, taking $v\in\mathbf{R}^{d}$ to be the nonzero vector given by
$v_{i}=\begin{cases}\beta_{i}&\text{if }i\in I,\\ -\beta_{i}&\text{if }i\in J,\\ 0&\text{otherwise,}\end{cases}$
we see that the tangent vector $\Delta=\beta v^{\top}+v\beta^{\top}$ is nonzero and (134) implies
$\nabla^{2}f(M_{\star})[\Delta,\Delta]=\sum_{i,j:\,\{i,j\}\in E^{*}_{\ell}}(v_{i}\beta_{j}+v_{j}\beta_{i})^{2}p_{ij}=\sum_{i,j:\,\{i,j\}\in E^{*}_{\ell}}(\beta_{i}\beta_{j}-\beta_{j}\beta_{i})^{2}p_{ij}=0,$ (135)
where the first equality holds since $v$ is supported on $V^{*}_{\ell}$. Thus, $\mu_{P}\in\mathcal{E}^{{\rm mc}}$, as claimed.
[…]
$\nabla^{2}f(M_{\star})[\Delta,\Delta]=\sum_{i,j:\,p_{ij}>0}(v_{i}\beta_{j}+v_{j}\beta_{i})^{2}p_{ij}=0$
since each summand vanishes. Therefore $\mu_{P}\in\mathcal{E}^{{\rm mc}}$, as claimed.
[…]
$v_{i}\beta_{j}+v_{j}\beta_{i}=0~~\text{whenever}~~\{i,j\}\in E.$ (136)
Now let $G_{\ell}^{*}=(V^{*}_{\ell},E^{*}_{\ell})$ be any connected component of $G^{*}$. Then $G^{*}_{\ell}$ is non-bipartite, so it must contain an odd-sized cycle $i_{1}\rightarrow i_{2}\rightarrow\dots\rightarrow i_{k}\rightarrow i_{1}$. If $k=1$, then $\{i_{1}\}\in E$ and hence (136) implies $2\beta_{i_{1}}v_{i_{1}}=0$. Similarly, if $k\geq 3$, then we have the expansion
$2\beta_{i_{1}}v_{i_{k}}=\big(\beta_{i_{1}}v_{i_{k}}+\beta_{i_{k}}v_{i_{1}}\big)-\frac{\beta_{i_{k}}}{\beta_{i_{2}}}\big(\beta_{i_{2}}v_{i_{1}}+\beta_{i_{1}}v_{i_{2}}\big)+\sum_{j=2}^{k-1}(-1)^{j}\frac{\beta_{i_{1}}\beta_{i_{k}}}{\beta_{i_{j}}\beta_{i_{j+1}}}\big(\beta_{i_{j+1}}v_{i_{j}}+\beta_{i_{j}}v_{i_{j+1}}\big)=0,$
[…]
@@
[…]
${\rm RSE}(\mu_{P})^{2}=\min_{\nu\in\mathcal{E}^{\text{mc}}}W_{2}^{2}(\mu_{P},\nu)$ $\leq W_{2}^{2}\big(\mu_{P},(\mathsf{P}_{A})_{\#}\mu_{P}\big)=\sum_{(i,j)\in\operatorname{supp}({P})\setminus A}p_{ij},$
where the last equality follows from Lemma A.1. Taking the minimum over all $A\in\Omega_{{\beta}_{\star}}$ such that $A\subset\operatorname{supp}(P)$ yields the inequality $\leq$ in (42).
Conversely, consider an arbitrary measure $\nu$ in $\mathcal{E}^{\text{mc}}$. Then $\nu=\mu_{Q}$ for some $Q\in\mathcal{Q}$ such that $\operatorname{supp}(Q)\in\Omega_{{\beta}_{\star}}$. Now set $B:=\operatorname{supp}(P)$ and $A^{\prime}:=\operatorname{supp}(Q)\cap B$; since $A^{\prime}$ is a symmetric subset of $\operatorname{supp}(Q)\in\Omega_{{\beta}_{\star}}$, it follows readily that $A^{\prime}\in\Omega_{{\beta}_{\star}}$. Moreover, clearly $A^{\prime}\subset B$ and the support of $(\mathsf{P}_{B})_{\#}\nu$ is contained in $\operatorname{range}(\mathsf{P}_{A^{\prime}})$, so (45) and Lemma A.1 imply
$W_{2}^{2}(\mu_{P},\nu)\geq W_{2}^{2}(\mu_{P},(\mathsf{P}_{B})_{\#}\nu)\geq W_{2}^{2}(\mu_{P},(\mathsf{P}_{A^{\prime}})_{\#}\mu_{P})=\sum_{(i,j)\in B\setminus A^{\prime}}p_{ij}\geq\min_{\begin{subarray}{c}A\in\Omega_{{\beta}_{\star}}\\ {A\subset B}\end{subarray}}\sum_{(i,j)\in B\setminus A}p_{ij},$
where the first and last inequalities hold trivially. Since $\nu\in\mathcal{E}^{\text{mc}}$ is arbitrary, this completes the proof of the reverse inequality $\geq$ in (42). Further, this argument shows that the $W_{2}$ distance from $\mu_{P}$ to $\mathcal{E}^{\text{mc}}$ is attained by the measure $(\mathsf{P}_{A^{*}})_{\#}\mu_{P}$ for any $A^{*}\subset\operatorname{supp}(P)$ in $\Omega_{{\beta}_{\star}}$ minimizing (42).
@@
Suppose that we have an algorithm Alg that outputs ${\rm RSE}(\mu_{P})^{2}$ for any given matrix completion data $(P,M_{\star})$, and let MaxCut be an algorithm that outputs the maximum cut size of any given undirected graph. We will construct a polynomial-time Turing reduction from MaxCut to Alg. To this end, let $G=(V,E)$ be an arbitrary undirected graph with connected components $G_{1}=(V_{1},E_{1}),\dots,G_{k}=(V_{k},E_{k})$. For each $\ell=1,\ldots,k$, let $d_{\ell}=|V_{\ell}|$, $V_{\ell}=\{n_{\ell}(1),\ldots,n_{\ell}(d_{\ell})\}$, and consider the instance of matrix completion data $(P_{\ell},M_{\star\ell})$ where $P_{\ell}\in\bm{S}^{d_{\ell}}$ is given by
$\big(P_{\ell}\big)_{ij}=\begin{cases}\frac{1}{2|E_{\ell}|}&\text{if }i\neq j\text{ and }\{n_{\ell}(i),n_{\ell}(j)\}\in E_{\ell},\\ \frac{1}{|E_{\ell}|}&\text{if }i=j\text{ and }\{n_{\ell}(i)\}\in E_{\ell},\\ 0&\text{otherwise}\end{cases}$
[…]
$|E_{\ell}|\cdot\texttt{Alg}({P}_{\ell},M_{\star\ell})=\min_{\begin{subarray}{c}A\in\Omega_{\beta_{\star\ell}}\\ {A\subset\operatorname{supp}(P_{\ell})}\end{subarray}}\sum_{\begin{subarray}{c}i\leq j\\ (i,j)\in\operatorname{supp}({P}_{\ell})\setminus A\end{subarray}}\mathbf{1}_{E_{\ell}}\big(\{n_{\ell}(i),n_{\ell}(j)\}\big)$
is equal to the minimum number of edges one needs to remove from $G_{\ell}$ so that the remaining subgraph is bipartite. Thus, $|E_{\ell}|\big(1-\texttt{Alg}({P}_{\ell},M_{\star\ell})\big)$ is equal to the number of edges of a maximum bipartite subgraph of $G_{\ell}$, which is equal to $\texttt{MaxCut}(G_{\ell}).$ Therefore summing over the connected components of $G$ yields
$\texttt{MaxCut}(G)=|E|-\sum_{\ell=1}^{k}|E_{\ell}|\cdot\texttt{Alg}({P}_{\ell},M_{\star\ell}),$
thereby providing a polynomial-time Turing reduction from the problem of computing MaxCut($G$) for any given undirected graph $G$ to the problem of computing ${\rm RSE}(\mu_{P})^{2}$ for any given matrix completion data $(P,M_{\star})$. Since the former problem is NP-hard in general, so is the latter.